Identify the mystery nucleus (represented by the question mark) in the following

reaction:

1 0 n + 235 92 U 144 56 Ba + ? + 2 1 0 n


1
Expert's answer
2021-03-01T07:19:54-0500

The reaction is given as : 0n1 + 92U235 = 56Ba144 + ? + 2 0n1

Solution :

let the mystery nucleus(?) be ZXA ; where Z is the atomic number and A is the mass number of the mystery nucleus

Now to identify the mystery nucleus(?) we have to equate both sum of the atomic number and the mass number for both sides of the given reaction

now equating sum of atomic number for both sides

0 + 92 = 56 + Z + 0

or, Z = 36

now equating sum of mass number for both sides

1 + 235 = 144 + A + 2

or, A = 90

then we can write the mystery nucleus as 36X90

now from the periodic table we know that the element having the atomic number 36 is known as Krypton

Answer :

The mystery nucleus(?) in the given reaction is given as 36Kr90